# XAI-CF – Examining the Role of Explainable Artificial Intelligence in Cyber Forensics

### v2

@@ 4.4.1 Summary
The reviewed papers collectively contribute significantly to the ever-evolving field of CF, a discipline vital for systematically collecting, analyzing, and preserving digital evidence to combat the rising tide of cybercrimes. Spanning various CF domains, the papers delve into sophisticated areas such as leveraging blockchain to enhance the trustworthiness of distributed file systems, addressing complex challenges in cloud forensics through provenance and cryptography, and refining log format unification for forensic soundness. These diverse contributions underscore the ongoing evolution of methodologies, techniques, and tools within the digital forensics landscape, emphasizing their pivotal role in ensuring the efficacy of cybercrime investigation and prevention strategies. The insights gleaned from these papers highlight the multifaceted nature of CF, where researchers grapple with intricate issues such as data integrity, privacy preservation, and the dynamic landscape of cyber threats. As the digital realm continues to expand, the need for innovative approaches to CF becomes increasingly evident. These papers collectively underscore the interdisciplinary efforts required to stay ahead of cybercriminals, pushing the boundaries of CF research and reinforcing its critical role in maintaining a secure and trustworthy digital environment.
@@
Types
Articles
General
(silde2014digital, ), (neasbitt2015webcapsule, ), (rahman2020new, ), (hina2021sefaced, ), (shahbazi2022nlp, ), (gupta2023digital, )
Network
(neasbitt2014clickminer, ), (nalawade2016forensic, ), (zhang2017blockchain, ), (nyaletey2019blockipfs, ), (dalezios2020digital, )
Mobile
(karpisek2015whatsapp, ), (anglano2016forensic, ), (luo2020automatic, ), (domingues2023post, )
IoT
(koroniotis2019towards, ), (noura2020distlog, ), (dragonas2023iot, )
Database
(kim2022methods, ), (nissan2023database, )
@@ 5 Explainable Artificial Intelligence in Cyber Forensics (XAI-CF)
Here we make a case for the significance and advantages of XAI in CF and strongly recommend its use in CF.
@@ 5.1 Why XAI-CF
The ability of an AI system to learn and adapt without requiring user input makes it beneficial in automating and completing different tasks in several fields of work. Also, its ability to deal with large amounts of data makes it ideal to be used in CF which deals with massive volumes of data (zawoad2015digital, ). Automation of evidence processing using AI-based techniques shows promising results in expediting the CF analysis process (du2020sok, ). Because of the inherent complexity of AI systems, one of the main problems with them is that they are not easy to understand and interpret, and it becomes difficult for humans to realize how an AI system came to a decision. This problem becomes more consequential when these systems are used in making important decisions, such as deciding a criminal case in a court of law where pieces of evidence are forensically analyzed and determined through an AI system. Here we raise an important question and issue of how to build AI models that will be interpretable and explainable not only to the experts but also to users without prior knowledge or expertise in the area of CF.
In general, the stakeholders of CF are not experts in AI. For them to make an informed decision, they need to first understand the output of the respective AI system. The use of XAI in CF improves the CF analysis and extracts forensically important pieces of evidence in a way that is easy to interpret and understand by the different stakeholders of CF, and hence can be used successfully in investigations and potentially in a court of law. We are already stepping into the digital age and if we want CF to take full advantage of this and successfully address the challenges (karie2015taxonomy, ; du2020sok, ) it faces, then we need to fully infuse AI into CF processes. To successfully infuse AI into CF we need to interpret and explain the results of an AI system to the stakeholders of CF. There are several studies (hepler2007object, ; keppens2012argument, ; fenton2013general, ; timmer2014extracting, ; verheij2014catch, ; vlek2014building, ; timmer2015structure, ; verheij2016arguments, ; vlek2016method, ; aditya2018enabling, ; afzaliseresht2019explainable, ; costantini2020assessing, ; bolle2020role, ; mahajan2021explainable, ; pethe2022atle2fc, ; hall2022proof, ; jayakumar2022visually, ; de2022interpretable, ; bouter2023protoexplorer, ) carried out to explain AI outputs for the stakeholders of CF involved in the decision-making process. We discuss some of them in detail in section 6.
@@ 5.2 How XAI-CF
When we talk about AI (AI-modern-approach-2020, ) two things come into mind, intelligence, such as the machines that perform functions that require intelligence when performed by people, and automation, such as decision-making, problem-solving, learning, etc. AI is a universal field (AI-modern-approach-2020, ) and encompasses a large variety of subfields, methods, and techniques, such as knowledge representation, adversarial searching, statistical and probabilistic reasoning, supervised, unsupervised and reinforcement learning, and natural language processing. In an AI system for CF, one or a combination of these techniques is used to perform CF analysis, and a decision (output) is reached. In an XAI system for CF, one or a combination of these techniques are also used to interpret and explain the decision (output) of the AI system. An XAI system can be divided into four categories (doran2017does, ), opaque, interpretable, comprehensible, and truly explainable, as shown in Table 4.
@@
Category
Description
Opaque
The mappings between input and output are invisible; a proprietary AI system where the licensor does not want to reveal the workings of the system; most difficult to interpret and explain.
Interpretable
The mappings between input and output are visible and understandable; a linear regression model where the feature weights can be compared to realize the relative importance of each feature; easier than opaque systems to interpret and explain.
Comprehensible
The system’s output is tagged with symbols (words or visualization) that allow the user to relate properties of the inputs to their output; the degree of comprehension about the symbols depends on the user’s knowledge about the AI system; a Bayesian network where edges are tagged with symbols to explain the transition; easier than interpretable systems to interpret and explain.
Truly Explainable
This is both interpretable and comprehensive; moreover, it provides reasoning that explains the decision-making process using human-understandable features of the input data; a simple decision tree where edges are tagged with symbols to explain the transition, and also provides formal and informal reasoning for each decision taken in the process; most easy to interpret and explain, and are human-understandable.
The first three categories encompass the current work on XAI-CF. The last category is very difficult to achieve in XAI-CF because of the critical nature of the CF processes and the inconsistent reasoning process of humans to understand a decision. A human-in-the-loop (nguyen2021human, ; solanke2022explainable, ) may help an XAI-CF model achieve true explainability. Here we discuss some of the techniques to achieve XAI-CF. Table 6 highlights these techniques in current and previous works on XAI-CF discussed in section 6.
@@ 5.2.1 Model Simplification
One of the basic approaches used by XAI models is to simplify the original model. A simplified model is easy to interpret, understand, and explain. There are different techniques to achieve this. One of the techniques used is a surrogate model (aditya2018enabling, ). Using the predictions of the original model a surrogate model is built, such as a simple decision tree. This transparent model is built when it is difficult to know the influence of the features on the prediction and is trained to approximate the behavior of the original model. The surrogate model helps understand the decision-making process of the original model. One of the problems with these surrogate models is that they may provide the same predictions as the original model but with different rationales and reasoning. This could produce contrasting relations among pieces of evidence that can lead forensics analysts and other stakeholders of CF to reach a different decision than without a surrogate model. To address this limitation, researchers are working (mariotti2023beyond, ) on developing metrics to evaluate faithful surrogate models in XAI. Another technique used is storytelling (afzaliseresht2019explainable, ) that highlights the semantic and inferred information from the model in a human-readable format. Other techniques (verheij2016arguments, ) used for formalizing and simplifying reasoning about the forensics legal evidence model are narrative, argumentative, and probabilistic approaches.
@@ 5.2.2 Feature Importance/Relevance
The most popular method used in XAI-CF is extracting the features, building the model, and then interpreting and explaining the model by computing the importance and relevance of the input features to the output of the model. Having this knowledge helps in crafting different inputs to the model for testing the robustness of the model and hence increases trust in the model. Feature importance helps in explaining the model using different techniques that reduce the efforts of humans to interpret the model. Feature relevance also helps in selecting the best model for forensics analysis. One such method (aditya2018enabling, ) first simplifies the model using a surrogate model and then computes the influence of the features on the prediction. (afzaliseresht2019explainable, ) mines temporal patterns from the model and uses storytelling to relate patterns to output and highlight the information in a human-readable format. (mahajan2021explainable, ) use LIME (ribeiro2016should, ) to relate the relevance of input features to the output and conclude that LIME is an important XAI technique to select the best model for forensics analysis.
@@ 5.2.3 Visualization
This technique in our opinion is the best approach to explain a model and is easy to understand by humans. It is highly desirable in an XAI-CF system to have an interactive, data-driven visualization tool for forensics analysis. This increases the transparency of a model and greatly improves the ability of a forensics analyst and other XAI-CF stakeholders to understand and reason about the model. This ability helps them in the final decision process. A saliency or a heat map visually explains the contributions of the model. Sometimes a graph of the model is tagged with symbols that relate the input to their output. Visualization of the output can greatly help in interpreting and explaining a model to a forensics analyst and other stakeholders of CF. One such work (keppens2012argument, ) extracts argument diagrams (tagged with symbols) visually representing a reasoning structure from the model. (bouter2023protoexplorer, ) visualize and interpret deep fake video data for forensics analysis. A forensics analyst can zoom in on a part of the image and observe the contributions of the model to the prediction score. As we can see there is a lot of room for improvement in this area and there is a need to develop new innovative methods to visualize CF models for better explanation, especially for non-experts. A human-in-the-loop can complement such a visualization to better understand and explain the model.
@@ 5.2.4 Human-In-The-Loop
A human-in-the-loop can complement an XAI-CF system to produce valuable, practical, and structured results. In our opinion, this is the most vital technique to achieve truly explainable systems. There are three approaches to including humans in AI: Active Learning (settles2009active, ) – where AI has more control than a human; Interactive Learning (amershi2014power, ) – where there is closer interaction between a human and AI; and Machine Teaching (ramos2020interactive, ) – where a human has more control than AI. For XAI-CF, we recommend using the middle approach the Interactive Learning. One of the main reasons for this recommendation is the nature of the data available in CF. Most of the data in CF is in binary form, such as hard disk raw data for file carving, malware programs, images, video, etc. This kind of data falls into the category of either semi-structured or unstructured data. To convert this raw data into a suitable internal representation for AI to detect or classify patterns, interactive learning is useful in giving additional structure to this raw data. Another main reason is the explainability that interactive learning can provide in an XAI-CF system. In addition to being part of the learning process in XAI-CF humans can be at the end interpreting what the model has learned. Adding forensics experts to the learning helps XAI-CF to improve its knowledge and in turn its explanation. Interactive learning also comes with some shortcomings (mosqueira2023human, ), such as the intermingling of AI and human aspects and the dependence on a human (forensics) expert. Currently, there are not enough works that study the role of human-in-the-loop in XAI (abdul2018trends, ). Recently in the area of XAI-CF, (nguyen2021human, ) proposed to use an expert forensics analyst as human-in-the-loop to provide adjustments so that with each iteration the system improves and the end produces a better explanation of the model.
@@ 5.3 Evaluating XAI-CF
There are metrics for evaluating an AI system and there is work in progress on developing metrics for evaluating the performance of a XAI system. How to measure the accuracy of an XAI system? There are certain criteria, such as, how good (precise and clear) is the explanation, how satisfied it is for the respective audience, and how trustful (justifiable) is the explanation. A detailed taxonomy of the XAI system’s evaluation methods is given in (lopes2022xai, ).
Most of the previous works propose metrics that only work with a specific XAI model (arras2022clevr, ; keane2021if, ; lakkaraju2017interpretable, ; nguyen2020quantitative, ; rosenfeld2021better, ; vilone2020comparative, ) or use humans to study and give their opinions about the usefulness of an XAI model (dieber2022novel, ; holzinger2020measuring, ). Only three approaches (hoffman2018metrics, ; rosenfeld2021better, ; sovrano2023objective, ) claim to be model agnostic and can be used with any XAI-CF model. Here we briefly discuss these three approaches.
Hoffman et al. (hoffman2018metrics, ) listed three major criteria to evaluate XAI; how precise and clear is the explanation; how satisfied it is for the respective audience; and how justifiable is the explanation. The authors recommend using researchers and experts to manually use these criteria to evaluate XAI models with the help of checklists, certain scales, and psychometrics.
Rosenfeld et al. (rosenfeld2021better, ) introduces four objective metrics to quantify XAI. Given the XAI goal, these metrics quantify the explanation itself and its appropriateness. The first metric compares the performance of the black-box model with a transparent model. XAI-CF demands a binary evaluation because of the involvement of legal issues. This metric can be used to evaluate the XAI-CF model, Any value greater than a threshold nullifies the explanation. The second metric measures the complexity of the application by the number of rules used in the explanation. The more simplifier an explanation is the more preferable it is. The third metric focuses on the number of features used to construct the explanation. The fourth metric measures the stability of the explanation. For all these metrics a threshold value is required which is computed by conducting user studies of the specific XAI model.
Recently, Sovrano et al. (sovrano2023objective, ) propose the degree (how much) of explainability as one of the metrics to measure the accuracy of the explanation described in a natural language. The authors use pre-trained deep language models for general-purpose answer retrieval (bowman2015large, ; reimers2019sentence, ; yang2019multilingual, ; karpukhin2020dense, ) to measure explainability. A graph is extracted by detecting, with a dependency parser, all the clauses within the explanation’s text. These clauses are represented as special triplets, subject, body, and object. A natural language representation is obtained of these triplets as a possible answer using pre-trained deep language models. The degree of explainability is then quantified by measuring its relevance to answering a pre-defined set of conventional questions.
@@ 5.4 Desiderata of XAI-CF
Here we discuss some of the main requirements and prerequisites of a successful and practical XAI-CF system. There are many studies (doran2017does, ; gunning2019darpa, ; phillips2021four, ; arrieta2020explainable, ; vilone2020explainable, ; lopes2022xai, ) that describe and discuss the prerequisites of a successful XAI system, such as Trustworthy, Transparent, Effective, Efficient, Interpretable, Informative, Clear, Simple, and so on. We take into consideration these prerequisites and add some of ours, and divide the requirements into four main categories and sixteen sub-categories as shown in Figure 8 that equip an XAI-CF model to become a successful and practical system.

### v3

@@ 4.4.1 Summary
The reviewed literature highlights notable advancements in digital forensics, focusing on improved methods for gathering evidence, tracking it, and analyzing it automatically. Many studies concentrate on specific applications like web browsers, encrypted messaging apps, mobile databases, social media, health-monitoring apps, and Android apps. These advancements help investigators recover hidden, deleted, or encrypted data and understand user activity. Another important area is managing data and ensuring log integrity. Many papers use blockchain and tracking methods to secure logging systems and maintain evidence custody, especially in cloud services and social media investigations. Memory forensics also helps detect malicious actions when logs are incomplete. The integration of machine learning, natural language processing (NLP), and deep learning supports tasks such as intrusion detection and email classification. Specific datasets and frameworks, like Bot-IoT and ForensicLLM, enhance automation and consistent analysis. Lastly, research on IoT connections and bot activity detection emphasizes the need for modern investigations to handle complex environments. Overall, these studies demonstrate progress toward more reliable, automated, and integrity-focused digital forensics, addressing new challenges in cloud systems, mobile apps, IoT networks, and cybercrime.
@@
Types Articles
General [123], [126], [127], [129], [135], [136]
Network [117], [119], [121], [122], [133]
Mobile [118], [125], [132], [134]
IoT [120], [124], [131]
Database [128], [130]
@@ 5 Explainable Artificial Intelligence in Cyber Forensics (XAI-CF)
Here, we make a case for the significance and advantages of XAI in CF and strongly recommend its use in CF.
@@ 5.1 Why XAI-CF
The ability of an AI system to learn and adapt without requiring user input makes it beneficial in automating and completing different tasks in several fields of work. Also, its ability to deal with large amounts of data makes it ideal to be used in CF, which deals with massive volumes of data [137]. Automation of evidence processing using AI-based techniques shows promising results in expediting the CF analysis process [4]. Because of the inherent complexity of AI systems, one of the main problems with them is that they are not easy to understand and interpret, and it becomes difficult for humans to realize how an AI system came to a decision. This problem becomes more consequential when these systems are used in making important decisions, such as deciding a criminal case in a court of law, where pieces of evidence are forensically analyzed and determined through an AI system. Here we raise an important question and issue of how to build AI models that will be interpretable and explainable not only to the experts but also to users without prior knowledge or expertise in the area of CF.
In general, the stakeholders of CF are not experts in AI. For them to make an informed decision, they need to first understand the output of the respective AI system. The use of XAI in CF improves the CF analysis and extracts forensically important pieces of evidence in a way that is easy to interpret and understand by the different stakeholders of CF, and hence can be used successfully in investigations and potentially in a court of law. We are already stepping into the digital age, and if we want CF to take full advantage of this and successfully address the challenges [138, 4] it faces, then we need to fully infuse AI into CF processes. To successfully infuse AI into CF, we need to interpret and explain the results of an AI system to the stakeholders of CF. There are several studies [139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157] carried out to explain AI outputs for the stakeholders of CF involved in the decision-making process. We discuss some of them in detail in section 6.
@@ 5.2 How XAI-CF
When we talk about AI [5], two things come to mind: intelligence, such as the machines that perform functions that require intelligence when performed by people, and automation, such as decision-making, problem-solving, learning, etc. AI is a universal field [5] and encompasses a large variety of subfields, methods, and techniques, such as knowledge representation, adversarial searching, statistical and probabilistic reasoning, supervised, unsupervised, and reinforcement learning, and natural language processing. In an AI system for CF, one or a combination of these techniques is used to perform CF analysis, and a decision (output) is reached. In an XAI system for CF, one or a combination of these techniques is also used to interpret and explain the decision (output) of the AI system. An XAI system can be divided into four categories [158], opaque, interpretable, comprehensible, and truly explainable, as shown in Table 4.
@@
Category Description
Opaque The mappings between input and output are invisible; a proprietary AI system where the licensor does not want to reveal the workings of the system; most difficult to interpret and explain.
Interpretable The mappings between input and output are visible and understandable; a linear regression model where the feature weights can be compared to realize the relative importance of each feature; easier than opaque systems to interpret and explain.
Comprehensible The system’s output is tagged with symbols (words or visualization) that allow the user to relate properties of the inputs to their output; the degree of comprehension about the symbols depends on the user’s knowledge about the AI system; a Bayesian network where edges are tagged with symbols to explain the transition; easier than interpretable systems to interpret and explain.
Truly Explainable This is both interpretable and comprehensive; moreover, it provides reasoning that explains the decision-making process using human-understandable features of the input data; a simple decision tree where edges are tagged with symbols to explain the transition, and also provides formal and informal reasoning for each decision taken in the process; most easy to interpret and explain, and are human-understandable.
The first three categories encompass the current work on XAI-CF. The last category is very difficult to achieve in XAI-CF because of the critical nature of the CF processes and the inconsistent reasoning process of humans to understand a decision. A human-in-the-loop [159, 33] may help an XAI-CF model achieve true explainability. Here, we discuss some of the techniques to achieve XAI-CF. Table 5 highlights these techniques in current and previous works on XAI-CF discussed in section 6.
@@ 5.2.1 Model Simplification
One of the basic approaches used by XAI models is to simplify the original model. A simplified model is easy to interpret, understand, and explain. There are different techniques to achieve this. One of the techniques used is a surrogate model [148]. Using the predictions of the original model, a surrogate model is built, such as a simple decision tree. This transparent model is built when it is difficult to know the influence of the features on the prediction, and is trained to approximate the behavior of the original model. The surrogate model helps understand the decision-making process of the original model. One of the problems with these surrogate models is that they may provide the same predictions as the original model but with different rationales and reasoning. This could produce contrasting relations among pieces of evidence that can lead forensic analysts and other stakeholders of CF to reach a different decision than without a surrogate model. To address this limitation, researchers are working [160] on developing metrics to evaluate faithful surrogate models in XAI. Another technique used is storytelling [149] that highlights the semantic and inferred information from the model in a human-readable format. Other techniques [146] used for formalizing and simplifying reasoning about the forensic legal evidence model are narrative, argumentative, and probabilistic approaches.
@@ 5.2.2 Feature Importance/Relevance
The most popular method used in XAI-CF is extracting the features, building the model, and then interpreting and explaining the model by computing the importance and relevance of the input features to the output of the model. Having this knowledge helps in crafting different inputs to the model for testing the robustness of the model and hence increases trust in the model. Feature importance helps in explaining the model using different techniques that reduce the efforts of humans to interpret the model. Feature relevance also helps in selecting the best model for forensics analysis. One such method [148] first simplifies the model using a surrogate model and then computes the influence of the features on the prediction. [149] mines temporal patterns from the model and uses storytelling to relate patterns to output and highlight the information in a human-readable format. [152] uses LIME [44] to relate the relevance of input features to the output and conclude that LIME is an important XAI technique to select the best model for forensics analysis.
@@ 5.2.3 Visualization
This technique, in our opinion, is the best approach to explain a model and is easy to understand by humans. It is highly desirable in an XAI-CF system to have an interactive, data-driven visualization tool for forensics analysis. This increases the transparency of a model and greatly improves the ability of a forensics analyst and other XAI-CF stakeholders to understand and reason about the model. This ability helps them in the final decision process. A saliency or a heat map visually explains the contributions of the model. Sometimes a graph of the model is tagged with symbols that relate the input to its output. Visualization of the output can greatly help in interpreting and explaining a model to a forensics analyst and other stakeholders of CF. One such work [140] extracts argument diagrams (tagged with symbols) visually representing a reasoning structure from the model. [157] visualize and interpret deep fake video data for forensics analysis. A forensics analyst can zoom in on a part of the image and observe the contributions of the model to the prediction score. As we can see, there is a lot of room for improvement in this area, and there is a need to develop new innovative methods to visualize CF models for better explanation, especially for non-experts. A human-in-the-loop can complement such a visualization to better understand and explain the model.
@@ 5.2.4 Human-In-The-Loop
A human-in-the-loop can complement an XAI-CF system to produce valuable, practical, and structured results. In our opinion, this is the most vital technique to achieve truly explainable systems. There are three approaches to including humans in AI: Active Learning [161] – where AI has more control than a human; Interactive Learning [162] – where there is closer interaction between a human and AI; and Machine Teaching [163] – where a human has more control than AI. For XAI-CF, we recommend using the middle approach the Interactive Learning. One of the main reasons for this recommendation is the nature of the data available in CF. Most of the data in CF is in binary form, such as hard disk raw data for file carving, malware programs, images, video, etc. This kind of data falls into the category of either semi-structured or unstructured data. To convert this raw data into a suitable internal representation for AI to detect or classify patterns, interactive learning is useful in giving additional structure to this raw data. Another main reason is the explainability that interactive learning can provide in an XAI-CF system. In addition to being part of the learning process in XAI-CF, humans can be at the end, interpreting what the model has learned. Adding forensics experts to the learning helps XAI-CF to improve its knowledge and, in turn, its explanation. Interactive learning also comes with some shortcomings [164], such as the intermingling of AI and human aspects and the dependence on a human (forensics) expert. Currently, there are not enough works that study the role of human-in-the-loop in XAI [165]. Recently, in the area of XAI-CF, [159] proposed to use an expert forensics analyst as a human-in-the-loop to provide adjustments so that with each iteration, the system improves and the result produces a better explanation of the model.
@@ 5.3 Evaluating XAI-CF
There are metrics for evaluating an AI system, and there is work in progress on developing metrics for evaluating the performance of an XAI system. How to measure the accuracy of an XAI system? There are certain criteria, such as how good (precise and clear) the explanation is, how satisfied it is for the respective audience is, and how trustworthy (justifiable) the explanation is. A detailed taxonomy of the XAI system’s evaluation methods is given in [69].
Most of the previous works propose metrics that only work with a specific XAI model [166, 167, 168, 169, 67, 170] or use humans to study and give their opinions about the usefulness of an XAI model [171, 172]. Only three approaches [66, 67, 68] claim to be model agnostic and can be used with any XAI-CF model. Here we briefly discuss these three approaches.
Hoffman et al. [66] listed three major criteria to evaluate XAI: how precise and clear the explanation is; how satisfying it is for the respective audience, and how justifiable the explanation is. The authors recommend using researchers and experts to manually use these criteria to evaluate XAI models with the help of checklists, certain scales, and psychometrics.
Rosenfeld et al. [67] introduce four objective metrics to quantify XAI. Given the XAI goal, these metrics quantify the explanation itself and its appropriateness. The first metric compares the performance of the black-box model with a transparent model. XAI-CF demands a binary evaluation because of the involvement of legal issues. This metric can be used to evaluate the XAI-CF model. Any value greater than a threshold nullifies the explanation. The second metric measures the complexity of the application by the number of rules used in the explanation. The more simplifier an explanation is, the more preferable it is. The third metric focuses on the number of features used to construct the explanation. The fourth metric measures the stability of the explanation. For all these metrics, a threshold value is required, which is computed by conducting user studies of the specific XAI model.
Recently, Sovrano et al. [68] propose the degree (how much) of explainability as one of the metrics to measure the accuracy of the explanation described in natural language. The authors use pre-trained deep language models for general-purpose answer retrieval [173, 174, 175, 176] to measure explainability. A graph is extracted by detecting, with a dependency parser, all the clauses within the explanation’s text. These clauses are represented as special triplets: subject, body, and object. A natural language representation is obtained of these triplets as a possible answer using pre-trained deep language models. The degree of explainability is then quantified by measuring its relevance to answering a pre-defined set of conventional questions.
@@ 5.4 Desiderata of XAI-CF
Here we discuss some of the main requirements and prerequisites of a successful and practical XAI-CF system. There are many studies [158, 177, 47, 26, 71, 69] that describe and discuss the prerequisites of a successful XAI system, such as Trustworthy, Transparent, Effective, Efficient, Interpretable, Informative, Clear, Simple, and so on. We take into consideration these prerequisites and add some of ours, and divide the requirements into four main categories and sixteen sub-categories as shown in Figure 9 that equip an XAI-CF model to become a successful and practical system.